a body is projected vertically upward with an unknown velocity.if acceleration due to gravity is 10m/s2 then distance travelled by the body in last second of upward motion is?

Grade:11

1 Answers

Chandan
6 years ago
Very good Question
in the last second means the ball will come to rest after one second
therefore using v+u+at
0=u-10t    (g=-10)(as opp to velocity)
as u is its initial velocity
for that second
therefore u=10m/s
use the formula for nth second distance
s(nth)=u+(a(2n-1)/2)
s=15 m
please dont forget to approve i need points urgently